OpenStudy (anonymous):

Can anybody tell me what 10b+9=29 please? :D xxx

OpenStudy (katrinakaif):

10b + 9 = 29 First, bring constants to one side. In this case, we'll have to substract 9 from both sides

OpenStudy (anonymous):

10b + 9 = 29 10b = 29-9 10b = 20 b = 20/10 b = 2

OpenStudy (katrinakaif):

10b + 9 = 29 - 9 - 9 10b = 20 Then divide by 10 both sides: 10b = 20 --- --- 10 10 b = 2